OpenStudy (anonymous):

MEDAL AWARDED someone please help :'( i'm stuck 2.2. A sample of radioactive material has been decaying for 5 years. Three years ago( at 2 years), there were 6.0 grams of material left. Now (at 5 years), there are 5.2 grams of material left. a. Write two point from this problem: _________ and _________ b. Using point-ratio form, y=y1b^x-x1, plug in the two points and solve for the ratio, b. Show all steps c.Write an exponential equation in point-ratio form for this situation D.Define your variables (include units)

OpenStudy (anonymous):

e. Use your equation to determine how much material will be left after 50 years from the beginning so x=50 f. what is the half-life of this material? That means when will there be half the original amount?

OpenStudy (anonymous):

Work that I've done a. two points are (2, 6.0) and (5,5.2) b.6.0=5.2^5-2 c? D? and from there i'm just lost on e and f

OpenStudy (wolf1728):

I figured I'd answer this because I know half life mathematics pretty well. However, I don't know it from a point of view that can express all the terms used in calculus. Here's a good link for half-life calculations: http://www.1728.org/halflife.htm 6 grams two years ago 5.2 grams now half-life = elapsed time * log (2)/ log (beginning amount/ending amount) half-life = 2 years * 0.3010299957 / log (6/5.2) half-life = 2 years * 0.3010299957 / 0.0621479067 half-life = 9.6875345096 years

OpenStudy (anonymous):

no haha this is just basic algebra 2 :)

OpenStudy (wolf1728):

e. Ending Amount = Beginning Amount / 2^n where 'n' is the elapsed time divided by half life Let's use the 2 year amount of 6 grams as the beginning time. Ending Amount = 6 g / 2^(50years) / 9.6875345096 years Ending Amount = 6 g / 2 ^ 5.1612719367 Ending Amount = 6 g / 35.7847237346 Ending Amount = 0.1676693118 grams

OpenStudy (wolf1728):

f. half-life has already been calculated 9.6875345096 years

OpenStudy (anonymous):

omgg thank you so much! From a-d i did a and b

OpenStudy (wolf1728):

okay - only trouble is I only know how to express this stuff in plain English instead of "two-point" or "point-ratio" terms.

OpenStudy (anonymous):

oh i see.. yea it's just referring back to the equation

OpenStudy (wolf1728):

Still, wouldn't it be better if they worded these problems so you could actually understand them?

OpenStudy (anonymous):

yea haha it's just a and b seem understandable from c and d it's just confusing

OpenStudy (wolf1728):

You already have done parts c and d?

OpenStudy (anonymous):

I have a guess for c. i think it's 5.2=6.0b^5-2

OpenStudy (wolf1728):

I could make a guess for part 'c' but I have NO idea about expressing it in "point-ratio" form.

OpenStudy (anonymous):

Well thank you so much you've done a whole lot already.. again thanks so much for your time :)

OpenStudy (wolf1728):

u r welcome - glad I saw your posting -I had a feeling I might be able to help.

OpenStudy (anonymous):

:))

OpenStudy (wolf1728):

okay see you later